### v2

@@ 3.2 Generator Variants for Part Assembly
Then we discuss different generator variants to model relationship between parts and predicts their corresponding 3D poses.
Parallel Generator. As illustrated in Figure 3 (c1), this is the the most straightforward way adopted by 3D-PA approaches [52, 53]. The parallel generator simultaneously outputs the 3D poses of the sequential parts in parallel, which can be denoted as:
$\small\{y_{i}\}_{i=1}^{N}=\mathcal{F}(\{p_{i}\}_{i=1}^{N})$ (3)
where $\mathcal{F}$ denotes the transformer encoder [43] with $L$ layers, which aims to build relationships among parts using self-attention.
Autoregressive Generator. Instead of parallel generation, another alternative is to autoregressively decode the 3D poses of object parts following the assembly sequence, where the generation progress is formulated as:
$\small y_{i}=\mathcal{F}(\{y_{j}\}_{j=1}^{i-1};\{p_{k}\}_{k=1}^{N}),i=1,2,\cdots,N$ (4)
To provide a global context of the potential assembled object, all part features are prepended as prompts before decoding the poses. At each step, the model generates a pose, concatenates it with the next part feature, and feeds it to the transformer model as input recursively. To decide the next part and maintain consistency in dimension, a part specifier indexes the upcoming part following the assembly sequence and projects the part feature for concatenation with the last pose feature, as depicted in Fig. 3 (c2). We use two linear projectors $\phi_{1}$ and $\phi_{2}$ to downsample the part and pose features from $v_{i}\in\mathbb{R}^{d}$ to $v_{i}\in\mathbb{R}^{\frac{d}{2}}$, where $d$ is the hidden dimension of the transformers.
@@ 3.3 Training Process
Training Loss. Supervision of both individual parts and the entire shape is applied during training. The part pose $y_{i}$ is represented by translation $t_{i}$ and rotation $r_{i}$, we supervise the translation via a MSE loss:
$\mathcal{L}_{t}=\sum\limits_{i=1}^{N}||t_{i}-t_{i}^{*}||_{2}^{2}$ (5)
@@
$\mathcal{CD}(\mathcal{X}_{1},\mathcal{X}_{2})=\sum\limits_{i,j\in\{1,2\}}\sum\limits_{x_{i}\in\mathcal{X}_{i}}\mathop{\min}_{x_{j}\in\mathcal{X}_{j}}||x_{i}-x_{j}||_{2}^{2}$ (6)
@@
$\mathcal{L}=\lambda_{t}\cdot\mathcal{L}_{t}+\lambda_{r}\cdot\mathcal{L}_{r}+\lambda_{s}\cdot\mathcal{L}_{s}$ (9)
@@ 4.1 Experiment Setup
Dataset Setup. We notice that prior studies only consider assembling merely three object categories (chair, table, lamp) in PartNet [32], which is a far cry from diverse objects in reality. To this end, we expand the available objects to include 21 categories, thereby introducing a more diversified set of objects to verify the SPAFormer’s generalization. We refer to this enhanced benchmark as PartNet-Assembly, and it is split in two subsets: (1) furniture object, containing three structurally complex categories (chair, table and storage) that offer sufficient and diverse samples, and (2) daily objects, containing the remaining 18 categories that represent commonly used everyday items (e.g. earphone). In total, PartNet-Assembly contains 22,873 shapes from 21 object categories with 173,765 distinct parts. The dataset is partitioned into training (70%), validation (10%) and testing (20%) sets, following the distribution protocal proposed by Mo et al. [32]. Furthermore, all target shapes within PartNet-Assembly are normalized to the canonical part space using PCA.
Implementation Details. We utilize the Adam optimizer initialized with a learning rate of 1.5e-4, and employ an exponentially decay strategy, reducing the learning rate by a factor of 0.8 every 80 epochs. The training process is conducted on 2 GPUs, employing a batch size of 128, over a total of 800 epochs. The optimal epoch is selected through performance assessment using Part Accuracy (PA) metrics on the unseen validation split. The transformer architecture is configured with a hidden dimension size of 512 across 6 layers. Following Huang et al. [13], the hyperparameters are set as follows: $\lambda_{t}=1$, $\lambda_{r}=10$, $\lambda_{s}=1$.
@@
    Shape Chamfer Distance (SCD) [25], which measures the overall distance between the predicted and ground truth shapes using the equation below, with all values scaled by $10^{3}$ for simplicity:
@@
    Part Accuracy (PA) [25], which assesses the correctness of part placement. A part is considered accurately placed if its error distance is below a specified threshold $\epsilon=0.01$:
    $\text{PA}=\mathcal{CD}(y_{i}(x_{i}),y_{i}^{*}(x_{i}))<\epsilon$ (11)
@@
    $\text{CA}=\frac{1}{|C|}\sum_{(c_{ij},c_{ji})\in C}((y_{i}(c_{ij})-y_{j}(c_{ji}))^{2}<\tau)$ (12)
@@
    Success Rate (SR) [26], which is defined as 1 for an object if the part accuracy for all its components equal to 1.
@@ 4.2 Ablation Study
In Table 4.2, we compare the performance of different variants outlined in Section 3.2 on the table assembly task. Both variants implement identical knowledge enhancement strategies. The results indicate that the parallel generator (denoted as "para") outperforms the autoregressive generator (denoted as "auto") across a majority of the 3D-PA metrics, with the exception of CA. More specifically, the "para" generator excels in SCD, PA and SR, highlighting its proficiency in assembling objects that are globally coherent. Conversely, the "auto" generator brings gains in CA, indicating its ability to ensure better connectivity among adjacent parts of assembled objects. If not specifically mentioned, we adopts the “para” generator in our following experiments.
Table 1: Comparison of generator variants on table assembly, where “para” and “auto” denote the parallel and autoregressive generator.
@@
We carry out ablations in Table 4.2 to evaluate the effectiveness of different assembly sequences: a) diagonal, which involves assembling an object from one corner diagonally to the opposite corner (e.g., assembly an chair from a front leg to the back); b) top-to-bottom where objects are assembled from the top down, assuming canonical poses; c) bottom-to-top, which is the reverse process of Top-to-Bottom, where assembly starts from the bottom; d) descending size, which is determined by the sizes of object components, calculated via Principal Component Analysis (PCA), starting with the largest. The outcomes, as detailed in the table, reveal that the diagonal pattern outperforms the others. This superior performance can be attributed to its ability to sequentially assemble symmetric components (such as the four legs of a chair), a feat that is less efficiently accomplished in other patterns like top-to-bottom. Based on these findings, we adopt the diagonal assembly pattern as our default strategy in subsequent experiments.
Table 3: Comparison of different knowledge enhancement strategies on chair assembly, where OEnc, REnc and SEnc denote order, relation and symmetry encoding.
@@
Table 3 reveals the impact of various encoding strategies on 3D part assembly. As can be seen, each strategy can bring improvements. For example, employing OEnc and SEnc augment the assembly outcomes by $+8.08\%$ and $+17.95\%$ on PA, by comparing Row 3-4 with Row 1. This demonstrates the value of incorporating order and symmetry information into the model, which significantly benefits the assembly process. However, while REnc alone does not notably advance assembly capabilities in Row 2, its combination with other encodings acts as a pivotal enhancer, akin to a catalyst. For instance, the comparison between Row 6 and Row 7 revels that the collaborative effect of REnc with other encodings greatly improve SR by $+2.35\%$ and PA by $+1.71\%$. Qualitative examples are provided in Figure 5(a) to present the gains of integrating each encoding strategies.
@@
We conduct an analysis of varied assembly lengths for the identification of assembly patterns and symmetry information in Figure 4. It is generally accepted that assembling objects composed of more parts presents a higher degree of difficulty. To this end, we categorize the testing set into two sets based on the number of parts: short-horizon assembly, which consists of fewer than 10 parts, and long-horizon assembly, encompassing those with more than 10 parts. For example, a chair usually comprises fewer than 10 parts including four legs, a back, and a seat. Compared with the vanilla attention baseline, our model (denoted as “+C”) not only exhibits a significant performance improvement of $+16.05\%$ in short-horizon assembly, but also shows an impressive gain of $+21.01\%$ in the more complex long-horizon assembly. This indicates a robust adaptability across assembly complexities.
Further analysis with SEnc (marked as “+S”) demonstrates the use of SEnc as a versatile technique that uniformly enhance PA performance across different assembly lengths. Such results indicate that our model is able to effectively captures complex structural information, marking a promising avenue for tackling challenging long-horizon tasks. In contrast, existing approaches [33, 13] struggle with long-horizon assembly, with a decrease on PA exceeding $-21.9\%$.
5. Effect of multi-task joint training.
Table 4: Comparison between class-specific and universal implementations of SPAFormer, evaluated on 21 object categories.
SCD↓ PA↑ CA↑ SR↑
category-specific 6.54 57.79 49.95 28.80
multi-task 6.37 59.78 53.74 29.40
An ideal assembly agent should be proficient in assembling a wide array of objects through a single, unified model.However, the exploration of such a universal assembly methodology remains underrepresented in existing works, and the category-specific approach is still widely-adopted. To address this gap, our work delves deeper into the feasibility of a general-purpose assembly approach. In addition to vanilla category-specific SPAFormer (denoted as “category-specific”) where each model is tailored to a specific category, we also develop a multi-task model (denoted as “multi-task”) that is trained across all accessible object categories. We compare the two implementations in Table 4.
The results reveal that the multi-task model excels in assembling objects from categories with both abundant (e.g., tables, storage furniture) and scarce (e.g., mugs with 120 samples, bowls with 130 samples) training data (Details can be found in supplementary). To be specific, the multi-task model presents a notable enhancement of $+1.99\%$ on PA, +3.79% on CA, and +0.6% on SR, highlighting the efficacy of a general-purpose approach to part assembly. Given these positive outcomes, we intend to further explore towards this direction, focusing on identifying which categories most benefit from shared assembly knowledge and controlling the generated categories in future works.
@@
Figure 5: Visualizations of: (a) assembly results when enhancing knowledge by adding new encoding pattern in a stepwise way; (b) common failure cases in object assembly, including rare shapes, long-horizon assembly and wrong orientation.
Figure 6: Quanlitative results and comparisons on daily object assembly. Sequence-conditioned model (RGL, ours) outperforms condition-free methods (Score, DGL). Our SPAFormer also achieves better assembly than RGL on most categories.
4.3 Comparison with State of the Arts
Comparison of Furniture Assembly. Assembling furniture encounters a notable challenge due to the complex and varied shapes of structured objects, which often troubles existing assembly methods. As shown in Table 5, our model significantly outperforms both previous score-based and graph-based models in the assembly of structured furniture objects across all evaluated metrics. For instance, when assembling unseen table models, our category-specific approach (denoted as “Ours-specific”) notably exceeds the performance of the leading prior model, RGL [33], achieving improvements of $+4.39\%$ on PA, $+7.71\%$ on CA, and $+7.07\%$ on SR. These results demonstrate the efficacy of our proposed SPAFormer in handling the furniture assembly tasks.
[…]
@@ 4.3 Comparison with State of the Arts
Table 6: Comparison of daily object assembly on 18 categories.
[…]
@@
Table 7: Comparison of conditional assembly on the chair task.
@@
Comparison of Daily Object Assembly. To assess our model’s performance in broader scenarios, we conduct evaluation on the collection of 18 everyday objects (e.g. earphone, bottle), showcasing its capability to generalize across a wider range of items. As illustrated in Table 7, our SPAFormer demonstrates a substantial advantage over condition-free methods, achieving an improvement of at least $+18.39\%$ on SR, and surpassing another sequence-conditioned RGL method by $5.12\%$ on SR. Despite condition-free approaches [3, 13] performs better on SCD, they significantly underperform across other metrics. We attribute this to the optimization collapse, where the value of SCD loss is low but the assembled shapes are less reasonable, as shown in Row 1-2 in Figure 6.

### v3

@@ 3.2 Generator Variants for Part Assembly
Then we discuss different generator variants to predict the 3D poses of object parts.
Parallel Generator. As illustrated in Figure 3 (c1), this is the most straightforward way adopted by 3D-PA approaches [15, 60]. The parallel generator simultaneously outputs the 3D poses of the sequential parts in parallel, which can be denoted as:
$\small\{y_{i}\}_{i=1}^{N}=\mathcal{G}(\mathcal{F}(\{v_{i}\}_{i=1}^{N}))$ (3)
where $\mathcal{F}$ denotes the transformer encoder [48] with $L$ layers, which aims to build relationships among parts using self-attention, and $\mathcal{G}$ is a shared MLP that maps part representations to poses.
Autoregressive Generator. Instead of parallel generation, another alternative is to autoregressively decode the 3D poses of object parts following the assembly sequence.
To provide a global context of the potential assembled object, all part features are prepended as prompts before decoding the poses. At each step $i$, the model generates a pose feature $z_{i}$, concatenates it with the next part feature $v_{i+1}$, and feeds it to the transformer model as input recursively. To decide the next part and maintain consistency in dimension, a part specifier indexes the upcoming part following the assembly sequence and projects the part feature for concatenation with the last pose feature, as depicted in Fig. 3 (c2). We use two linear projectors $\phi_{1}$ and $\phi_{2}$ to downsample the part $v$ and pose $z$ features from $\mathbb{R}^{d}$ to $\mathbb{R}^{\frac{d}{2}}$, respectively, where $d$ is the hidden dimension of the transformers. As a result, the generation progress is formulated as:
$\small y_{i}=\mathcal{G}(\mathcal{F}(\{v_{k}\}_{k=1}^{N};\{(\phi_{1}(z_{j})\oplus\phi_{2}(v_{j+1})\}_{j=0}^{i-1}))$ (4)
@@ 3.3 Training Process
Training Loss. Supervision of both individual parts and the entire shape is applied during training. The part pose $y_{i}$ is represented by translation $t_{i}\in\mathbb{R}^{3}$ and rotation $r_{i}\in\mathbb{R}^{4}$ in the form of quaternion. The translation is supervised via a MSE loss:
$\mathcal{L}_{t}=\sum\limits_{i=1}^{N}\|t_{i}-t_{i}^{*}\|_{2}^{2}$ (5)
@@
$\mathcal{CD}(\mathcal{X}_{1},\mathcal{X}_{2})=\sum\limits_{i,j\in\{1,2\}}\sum\limits_{x_{i}\in\mathcal{X}_{i}}\min_{x_{j}\in\mathcal{X}_{j}}\|x_{i}-x_{j}\|_{2}^{2}$ (6)
@@
$\mathcal{L}=\lambda_{t}\mathcal{L}_{t}+\lambda_{r}\mathcal{L}_{r}+\lambda_{s}\mathcal{L}_{s}$ (9)
@@ 4.1 Experiment Setup
Dataset Setup. We notice that prior studies only consider assembling merely three object categories (chair, table, lamp), which is a far cry from diverse objects in reality. To this end, we expand the available objects to include 21 categories from PartNet [36], thereby introducing a more diversified set of objects to verify the SPAFormer’s generalization. We refer to this enhanced benchmark as PartNet-Assembly, and it is split into two subsets: (1) furniture object, containing three structurally complex categories (chair, table, and storage) that offer sufficient and diverse samples, and (2) daily object, containing the remaining 18 categories that represent commonly used everyday items (e.g. earphone). In total, PartNet-Assembly contains 22,873 shapes from 21 object categories with 173,765 distinct parts. The dataset is partitioned into training (70%), validation (10%), and testing (20%) sets, following the distribution protocol proposed by Mo et al. [36].
Implementation Details. We utilize the Adam optimizer initialized with a learning rate of 1.5e-4, and employ an exponential decay strategy, reducing the learning rate by a factor of 0.8 every 80 epochs. The training process is conducted on 2 GPUs, employing a batch size of 128, over a total of 800 epochs. The optimal epoch is selected through performance assessment using Part Accuracy (PA) metrics on the unseen validation split. The transformer architecture is configured with a hidden dimension size of 512 across 6 layers. Following Huang et al. [15], the hyperparameters are set as follows: $\lambda_{t}=1$, $\lambda_{r}=10$, $\lambda_{s}=1$. Furthermore, all target shapes and input parts within PartNet-Assembly are normalized to the canonical space using PCA, which eliminates the impact of initial poses.
@@
    Shape Chamfer Distance (SCD) [27], measures the overall distance between the predicted and ground-truth shapes using the equation below, with all values scaled by $10^{3}$ for simplicity:
@@
    Part Accuracy (PA) [27], assesses the correctness of part placement. A part is considered accurately placed if its error distance is below a specified threshold $\epsilon=0.01$:
    $\text{PA}=\mathbb{I}\left[\mathcal{CD}(y_{i}(x_{i}),y_{i}^{*}(x_{i}))<\epsilon\right]$ (11)
@@
    $\text{CA}=\frac{1}{|C|}\sum_{(c_{ij},c_{ji})\in C}\mathbb{I}\left[(y_{i}(c_{ij})-y_{j}(c_{ji}))^{2}<\tau\right]$ (12)
@@
    Success Rate (SR) [28], is defined as 1 for an object if the part accuracies of all components are equal to 1.
@@ 4.2 Ablation Study
In Table 1, we compare the performance of different variants outlined in Section 3.2 on the table assembly task. Both variants implement identical knowledge enhancement strategies. The results indicate that the parallel generator (denoted as ”para”) outperforms the autoregressive generator (denoted as ”auto”) across a majority of the 3D-PA metrics, with the exception of CA. Specifically, the ”para” generator excels in SCD, PA, and SR, highlighting its proficiency in assembling objects that are globally coherent. Conversely, the ”auto” generator brings gains in CA, indicating its ability to ensure better connectivity among adjacent parts of the objects. Due to the ease of training, we adopt the “para” generator in our following experiments.
Table 1: Comparison of different generator variants on table assembly, where “para” and “auto” denote the parallel and autoregressive generator, respectively.
@@
We carry out ablations in Table 2 to evaluate the effectiveness of different assembly sequences: a) diagonal, which involves assembling an object from one corner diagonally to the opposite corner (e.g., assembly a chair from a front leg to the back); b) top-to-bottom, where objects are assembled from the top down; c) bottom-to-top, which is the reverse process of top-to-bottom, where assembly starts from the bottom; d) descending size, which is determined by the sizes of object components, calculated via principal component analysis (PCA), starting with the largest. The outcomes reveal that the diagonal pattern outperforms the others. This superior performance can be attributed to its ability to sequentially assemble symmetric components (such as the four legs of a chair), a feat that is more efficiently accomplished than other patterns like top-to-bottom. Based on these findings, we adopt the diagonal assembly pattern as our default strategy in subsequent experiments.
Table 3: Comparison of different knowledge enhancement strategies on chair assembly, where OEnc, REnc, and SEnc denote order, relation, and symmetry encoding, respectively.
@@
Table 3 reveals the impact of various encoding strategies on 3D part assembly. As can be seen, each strategy can bring improvements. For example, employing OEnc and SEnc augment the assembly outcomes by $+8.08\%$ and $+17.95\%$ on PA, by comparing Row 3-4 with Row 1. This demonstrates the value of incorporating order and symmetry information into the model, which significantly benefit the assembly process. However, while REnc alone does not notably advance assembly capabilities in Row 2, its combination with other encodings acts as a pivotal enhancer, akin to a catalyst. For instance, the comparison between Row 6 and Row 7 reveals that the collaborative effect of REnc with other encodings improves SR by $+2.35\%$ and PA by $+1.71\%$. Qualitative examples are provided in Figure 4 to present the gains of integrating each encoding strategy.
@@
Figure 4: Visualizations of assembly results when enhancing knowledge by adding new encoding patterns in a stepwise way.
@@
Table 5: Comparison of daily object assembly on 18 categories.
@@
We conduct an analysis of varied assembly lengths for the identification of assembly patterns and symmetry information in Figure 5. It is generally accepted that assembling objects composed of more parts presents a higher degree of difficulty. To this end, we categorize the testing set into two sets based on the number of parts: short-horizon assembly, which consists of fewer than or equal to 10 parts, and long-horizon assembly, encompassing those with more than 10 parts. For example, a chair usually comprises fewer than 10 parts including four legs, a back, and a seat. Compared with the vanilla attention baseline, our model that captures sequence information (denoted as “+C”) not only exhibits a significant performance improvement of $+16.05\%$ in short-horizon assembly but also shows an impressive gain of $+21.01\%$ in the more complex long-horizon assembly. Further analysis with SEnc (marked as “+S”) demonstrates the use of SEnc as a versatile technique that uniformly enhances PA performance across different assembly lengths. Such results indicate that our model is able to effectively capture complex structural information, marking a promising avenue for tackling challenging long-horizon tasks. In contrast, existing approaches [37, 15] struggle with long-horizon assembly, with a decrease in PA exceeding $-21.9\%$.
5. Effect of multi-task joint training.
Table 6: Comparison between category-specific and universal implementations of SPAFormer, evaluated on 21 object categories.
SCD↓ PA↑ CA↑ SR↑
ours-specific 6.54 57.79 49.95 28.80
ours-multitask 6.37 59.78 53.74 29.40
An ideal assembly agent should be proficient in assembling a wide array of objects through a unified model. However, the exploration of such a universal assembly methodology remains underrepresented in existing works, and the category-specific approach is still widely adopted. To address this gap, our work delves deeper into the feasibility of a general-purpose assembly approach. In addition to vanilla category-specific SPAFormer (denoted as “ours-specific”) where each model is tailored to a specific category, we develop a multi-task model (denoted as “ours-multitask”) that is trained across all accessible object categories. We compare the two implementations in Table 6.
The results reveal that the multi-task model excels in assembling objects from categories with both abundant (e.g., tables, storage furniture) and scarce (e.g., mugs with 120 samples, bowls with 130 samples) training data (Details can be found in supplementary). To be specific, the multi-task model presents a modest enhancement of $+1.99\%$ on PA, +3.79% on CA, and +0.6% on SR, highlighting the viability of a general-purpose approach to part assembly. Given these positive outcomes, we intend to further explore towards this direction, focusing on identifying which categories most benefit from shared assembly knowledge and controlling the generated categories in future works.
4.3 Comparison with State of the Arts
Comparison of Furniture Assembly. Assembling furniture is challenging due to the complex and varied shapes of structured objects, which troubles many assembly methods. As shown in Table 4, our model significantly outperforms both previous score-based and graph-based models in the assembly of structured furniture objects across almost all evaluated metrics. For instance, when assembling unseen tables, our category-specific approach (denoted as “Ours-specific”) remarkably exceeds the performance of the leading prior model, RGL [37], achieving improvements of $+4.39\%$ on PA, $+7.71\%$ on CA, and $+7.07\%$ on SR. These results demonstrate the efficacy of our proposed SPAFormer in handling furniture assembly tasks.
Table 7: Comparison of conditional assembly on chair assembly.
@@ 4.3 Comparison with State of the Arts
Comparison of Daily Object Assembly. To assess our model’s performance in broader scenarios, we conduct experiments on the collection of 18 everyday objects (e.g. earphone, bottle), showcasing its capability to generalize across a wider range of items. As illustrated in Table 5, our SPAFormer demonstrates a substantial advantage over condition-free methods, achieving an improvement of at least $+18.39\%$ on SR, and surpassing another sequence-conditioned RGL method by $5.12\%$ on SR. Despite condition-free approaches [3, 15] performs better on SCD, they significantly underperform across other metrics. We attribute this to the optimization collapse, where the value of SCD loss is low but the assembled shapes are less reasonable, as shown in Row 1-2 in Figure 1.
4.4 Qualitative Analysis
Figure 6: Visualizations of failed object assemblies, including rare shapes, long-horizon assembly, and wrong orientation.
Failure cases. In exploring the limitations of our model, Figure 6 highlights two primary factors leading to assembly failures: out-of-distribution (OOD) shapes and our model’s inherent weakness. The left two columns in Figure 6 depict OOD shapes that are rarely encountered in reality, such as recliner chair and sofa without chair legs. The examples on the right side of Figure 6 shed light on inherent weaknesses in our model. Despite making significant strides in addressing long-horizon assembly challenges as evidenced in Figure 5, our model exhibits limitations in assembling complex objects, such as chairs composed of up to 17 parts. Furthermore, our model encounters difficulties in accurately recognizing part orientations, primarily due to its limited ability to detect the junctions between parts. To address these issues, labeling parts with peg-hole joints [29] and reframing the problem as bipartite matching is a costly alternative.
Real-world experiment. Additionally, we conduct a real-world experiment, where the object scan is sourced from the Redwood dataset [5], to successfully assemble a table by our model, suggesting potential generalization to real-world scenarios where the part point clouds are imperfect.
Figure 7: Real-world experiment on table assembly.
5 Conclusion
In this study, we introduce a novel model SPAFormer, which is designed to tackle the issue of combinatorial explosion in 3D part assembly tasks. Building upon insights from prior research, SPAFormer employs assembly sequences to effectively narrow down the extensive solution space. We explore different assembly generator variants and incorporate multiple knowledge enhancement strategies for capturing the inherent assembly pattern. To facilitate comprehensive evaluation, we extend the existing benchmark to PartNet-Assembly, which spans 21 diverse categories including both furniture and daily objects. Experimental analyses are conducted to validate components of SPAFormer and the enhanced performance over existing methods.
Acknowledgements
This work was partially supported by the Beijing Natural Science Foundation (No. L233008).
Supplementary Material
Appendix A Additional Quantitative Results
A.1 General-purpose Part Assembly is Beneficial
We have trained a versatile model (“ours-general”) on all available object categories, which aims to perform part assembly across a variety of categories, and category-specific models (“ours-specific”), where each model is tailored to a specific category.
Influence on Assembly Sequence Length. We analyze the impact of multi-task learning on varied assembly lengths in Tab. 8. Compared to the category-specific models, the multi-task model continues to gain a notable improvement of $+2.77\%$ in long-horizon assembly, and also improve by $+1.9\%$ in short-horizon assembly.
Table 8: Average results across all categories. The numbers in brackets denote the object number in the test set.
$\leq$10 parts (3318) $>$10 parts (1223)
Method PA SR PA SR
ours-specific 61.74 36.20 53.25 7.59
ours-multitask 63.64 37.46 56.02 9.31
Detailed Results on Daily Object Assembly. In  Tab. 9, we conduct comparisons between the versatile model and category-specific models. Our versatile model demonstrates significant improvements in categories with both abundant (e.g., tables, storage furniture) and limited (e.g., mugs with 120 samples, bowls with 130 samples) training data. Given these promising results, we will further explore this direction, such as discovering which categories benefit most from shared assembly knowledge, or how to effectively manage the output categories.
A.2 More Details and Analysis on Category-specialized Part Assembly
Unlike the less-explored general-purpose assembly mentioned above, the experiments presented in our main paper follow the traditional benchmark of category-specialized part assembly, which has been commonly adopted by previous works. In this subsection, we provide additional experimental details.
Additional Implementation Details. In our approach, each part is represented by a point cloud, consisting of 1,000 points obtained through Furthest Point Sampling [8]. We ensure these sampled point clouds are zero-centered and aligned with the principal axes determined by Principal Component Analysis (PCA) [18]. In our experimental results, the shape chamfer distance (SCD) metric is scaled up by a factor of 1000, while the part accuracy (PA), connectivity accuracy (CA), and success rate (SR) are expressed as percentages. Moreover, the assembly sequences are constructed based on the relative positions of parts in their original configuration.
Additional evaluation on Per-class Part Assembly. To offer a comprehensive comparison of different object categories, we have included per-class evaluation results in Tab. 9. Our proposed method consistently outperforms previous works in most categories. It’s important to note that the success rates vary among categories, We attribute this variation more to the diversity of object structures and fundamental geometries encountered during training, rather than the quantity of training samples. For example, categories with a large number of training samples, such as faucet (with approximately 460 samples) and lamp (with approximately 1420 samples), do not necessarily yield higher success rates. Specifically, faucets and lamps demonstrate success rates of only 5.3% and 17.2%, respectively, when using our category-specific model. This is largely due to the complex assembly required by the intricate geometrical composition of their parts.
Table 9: Comparisons of per-class results across 18 daily objects, where “ours-multitask” and “ours-specific” denote the unified model for multi-task assembly and the class-specific models, respectively. The abbreviations “Dish”, “Disp”, “Ear”, “Fauc” and “Frid” denote Dishwasher, Display, Earphone, Faucet and Refrigerator, respectively.
Method Avg Bag Bed Bottle Bowl Clock Dish Disp Door Ear Fauc Hat Knife Lamp Lap Mug Frid Trash Vase
DGL 6.57 23.7 20.2 6.0 20.2 3.1 7.0 1.4 4.5 16.6 13.8 3.6 2.0 8.0 0.9 5.2 5.8 9.1 2.4
RGL 10.93 11.8 9.6 3.5 22.4 9.1 4.5 3.6 3.7 27.2 17.2 9.6 3.4 15.8 17.1 7.8 4.2 4.7 8.7
Score 6.38 1.3 19.8 13.6 0.6 7.1 7.8 2.9 9.7 13.6 9.0 1.6 10.4 12.7 12.7 1.4 2.5 3.6 1.2
ours-specific 8.61 7.6 7.7 3.0 23.3 7.7 4.3 3.3 3.4 24.5 11.2 9.2 3.3 11.7 5.1 8.5 2.8 4.7 8.9
SCD ours-multitask 8.60 8.5 12.0 3.6 10.7 8.8 3.5 4.4 6.5 24.4 12.4 10.8 3.3 11.6 4.0 7.0 2.9 2.7 8.7
DGL 36.55 21.7 11.8 80.4 63.0 45.2 11.3 62.9 26.7 27.8 13.1 36.7 41.0 32.8 65.2 13.1 16.5 11.2 23.3
RGL 51.33 36.2 13.3 86.1 46.3 52.0 45.2 82.6 56.7 41.0 33.7 55.1 82.0 33.3 60.0 59.5 48.5 49.4 47.3
Score 32.39 62.8 1.5 2.2 72.2 11.6 1.8 10.2 10.2 14.0 23.5 58.2 59.8 34.9 62.8 54.8 32.5 20.6 51.4
ours-specific 54.32 52.2 28.0 87.8 42.6 58.4 52.5 83.6 47.1 52.0 39.7 62.2 76.2 34.2 87.2 37.0 52.1 59.4 48.5
PA ours-multitask 59.93 46.4 11.0 84.8 77.8 56.0 62.0 82.2 42.0 50.5 42.8 55.1 84.0 36.9 90.2 55.9 55.7 65.6 52.7
DGL 40.01 29.8 21.8 76.5 64.3 41.1 14.7 66.3 58.1 38.0 29.7 36.5 42.5 38.7 31.3 9.4 28.7 10.8 27.4
RGL 50.52 23.4 12.6 76.5 71.4 36.9 27.5 76.0 83.8 42.2 47.5 50.0 74.0 47.4 71.3 11.3 37.3 20.7 32.21
Score 32.16 16.2 19.4 24.8 50 26.2 10.9 19.8 19.6 24.1 37.8 55.8 51.4 41.4 16.2 11.3 32.5 9.8 38.5
ours-specific 51.60 42.6 23.2 77.1 57.1 37.6 29.4 80.2 70.4 55.1 44.2 51.9 77.3 46.9 95 1.9 50.4 27.6 26.4
CA ours-multitask 52.86 42.6 15.0 71.2 42.8 41.1 56.8 77.8 63.7 55.1 53.9 44.2 76.2 47.2 91.2 9.4 55.6 33.0 31.2
DGL 18.73 13.8 0 69.0 53.8 23.9 0 37.7 3.9 0 0 13.3 7.8 14.5 47.6 0 0 0 11.6
RGL 32.0 17.2 0 75.0 25.6 47.7 13.7 67.5 47.1 1.9 1.5 24.4 59.7 14.0 41.5 31.4 12.9 26.3 40.0
Score 15.05 34.5 0 0 66.7 3.4 0 8.4 11.8 0 1.5 26.7 19.5 10.1 12.2 25.7 3.2 2.6 43.6
ours-specific 37.12 31.0 6.3 78.6 25.6 53.4 15.7 69.7 45.1 17.0 5.3 37.8 50.6 17.2 78.0 5.7 16.2 42.1 42.2
SR ours-multitask 39.06 24.1 0 75.0 71.8 46.6 25.5 68.0 33.3 17.0 10.6 26.7 62.3 15.7 82.9 31.4 19.4 39.5 47.8
Test Number 29 16 84 39 88 51 191 51 53 132 45 77 407 82 35 31 38 232
Appendix B Additional Qualitative Examples
Additional visualizations, including both successful and unsuccessful cases, are presented in Figs. 8, 9 and 10. Our approach not only excels in effectively assembling structural tables that have small connection areas but also shows significant improvement in assembling larger parts of storage furniture, which typically have more extensive connection areas.
Appendix C Limitations and Directions
First, acquiring the assembly sequence is sometimes unfeasible, especially when predicted by assembly sequence planning algorithms. On the other hand, while assembly sequences can be derived from various sources such as instructional manuals, the goal for a truly autonomous agent is to independently determine assembly sequences and navigate the assembly process on its own, mirroring human capabilities. Therefore, combining the task of assembly sequence planning and 3D part assembly would significantly benefit the object assembly problem.
Second, current methods for 3D-PA primarily focus on point cloud processing without motion planning in the real world, which limits the application in robotics. Working towards this path and deploying the models in the real world can significantly expand the scope of this problem.